Posie’s plush puppy collection is now 24. Her collection is growing at a rate of about 50% per year. Copy and complete the table to predict the population for the next 5 years. Round to the appropriate place value.

Years 0 1 2 3 4 5
Population 24
Write a model. Graph the model on your calculator and find how long it will take to reach 1384 puppies? Round to the nearest tenth.

Answer by jorel1380(3719)

Year 0:24
Year 1: 24+12=36
Year 2: 36+18=54
Year 3: 54+27=81
Year 4: 81+40.5=121.5
Year 5: 121.5+60.75=182.25
If her collection is growing by 50% a year, then
1384=24 x (1+.5)^t.....where t is years. So:
1384/24=1.5^t
ln 57.6667=ln 1.5^t
ln 57.6667=t ln 1.5
t=ln 57.6667/ln 1.5=10
In approximately 10 years she will have 1384 plush puppies. ☺☺☺☺